# HLSFactory: A Framework Empowering High-Level Synthesis Datasets for Machine Learning and Beyond

### v2

@@ Abstract
In addition to its evident application in ML, we showcase the versatility and multi-functionality of our framework through six case studies: I) Using design space sampling in stage 1 to expand the design space covered form a small base set of HLS designs; II) Demonstrating the speedup from the fine-grained design parallelism backend; III) Extending HLSFactory to target Intel’s HLS flow across all stages; IV) Adding and running new auxiliary designs using HLSFactory; V) Integration of previously published HLS data in stage 3; VI) Using HLSFactory to perform HLS tool version regression benchmarking to identify differences between two versions of the same HLS vendor tool.
@@
I Introduction
@@
Machine learning (ML) techniques have been widely applied to different electronic design automation (EDA) flows including high-level synthesis (HLS)  [1, 2, 3, 4, 5, 6, 7, 8, 9, 10] for quality-of-result (QoR) prediction, optimization, and design space exploration (DSE). A key enabler to the success of such ML techniques is high-quality datasets, most of which are developed for individual studies e.g., [1, 9, 8]. Some recent works have contributed open-source datasets that can be used by other reseachers, e.g.,  [11, 12, 13, 14, 15, 16].
@@
Despite the great benefits of these datasets, there are still fundamental limitations that hinder their wider adoption for ML applications and FPGA research. First, these datasets are usually small or homogeneous, containing only a subset of previously published HLS benchmarks [17, 18, 19, 12], and frequently consisting exclusively of designs that work with one HLS tool from a single vendor. For example, Spector [11] contains only 9 Intel HLS designs, HLSDataset [13] contains 34 AMD/Xilinx HLS designs, and Rosetta [12] contains 6 AMD/Xilinx HLS designs. Second, because of these separately developed HLS datasets, the designs and intermediate/final tool outputs, which serve as important ML model features, are often reported organized in non-standard ad hoc ways. Some datasets contain only source code [18, 17, 19], some datasets contain only resource usage and end-to-end throughput [12] but no clock frequency or power numbers, while some contain only post-implementation results [16]. HLSyn [16] is a dataset for HLS designs targeted towards predicting design quality of FPGAs. It consists of a wider range of programs and compiler directives, enabling performance optimization of designs. However, existing datasets require huge manual effort and deep domain-specific knowledge for ML practitioners if they need a complete, unified, and larger dataset, where they must execute all related HLS tools on their own to re-collect and organize the needed information. Third, it is challenging for external users who want to extend the existing datasets by contributing their own designs, primarily caused by ad-hoc data formats and missing details when building these datasets (e.g., tool version, target FPGA device, clock frequency, implementation flow settings). The fundamental limitation is, however, not the lack of another complete and rich HLS dataset, but rather the lack of a flexible and extensible framework to enable continuous contributions to a standardized and sustainable dataset.
@@
  • Multi-purpose. Beyond simply being used to build datasets for ML training, HLSFactory is useful for any task where a large, diverse set of HLS runs is needed. This can include regression testing between various HLS tool vendors and versions or benchmarking published academic HLS tools and techniques. Users are able to explore different research domains quickly due to the aforementioned extensibility provided by our work allowing them to carry out experiments with custom flows and tools.
  • High performance and open-source. HLSFactory is capable of running tool flows, such as HLS synthesis, with a user-tunable, fine-grained multiprocessing backend that allows for automatic parallelism across multiple source datasets of HLS designs. We release HLSFactory as an open-source project available on GitHub. HLSFactory provides an end-to-end framework and an initial set of built-in HLS design sources mentioned in this work. It is released as a Python package (with design sources built into the package installation) along with documentation and demos to help users get started with minimal hassle.
In Sec. II, we first provide background on the prior works in existing HLS benchmarks and datasets. Sec. III introduces our HLSFactory framework detailing the three stages. Sec. IV dives into the implementation of HLSFactory, including how it is configured and extended, and our fine-grained parallelism technique to speed up dataset generation. We then perform several case studies in Sec. V that demonstrate the multi-purpose of the proposed framework.
TABLE I: A comparison of HLSFactory with the existing work. : feature supported;
: feature unsupported; : feature partially supported.
@@
II Related Work
HLS community has multiple standard benchmarks for assessing HLS tools including PolyBench [17], CHStone [18], and MachSuite [19], which in total provide around 67 benchmark designs and are far from sufficient for ML training. Rosetta [12], Dai [2], MLSBench [15], DB4HLS [14], HLSDataset [13], and Spector [11] are all recently proposed HLS datasets, where the former four use AMD/Xilinx tools and the last uses Intel tools. MLSBench provides a sampling from different combinations of directives (pragmas) on top of CHStone and MachSuite. DB4HLS provides exhaustive design exploration on 39 designs from MachSuite with a domain-specific language (DSL) for DSE and parallelized synthesis runs. HLSDataset aims to cover all four commonly used benchmarks (PolyBench, CHStone, MachSuite, Rosetta) with a DSL for specifying the design space to sample from. They also illustrate two ML-based case studies for post-implementation resource and power prediction. HLSyn [16] uses control data flow graphs (CDFGs) of compiled HLS kernels for QoR prediction using graph neural network approaches; their designs are sampled from PolyBench and MachSuite. The features of selected prior works and HLSFactory are shown in Table I.
@@ II Related Work
III HLSFactory Framework
Figure 1: A complete overview of the HLSFactory framework with three stages and three entry points where users can contribute their own designs.
III-A HLSFactory Overview
@@
Stage ➊ is design space expansion, aiming at expanding a single HLS design into multiple designs by enumerating different combinations of optimization directives (pragmas), which can significantly increase the number of data points for ML. In this stage, users can submit one or more HLS designs with possible DSE configurations and HLSFactory will extrapolate and expand the design space via a frontend. This frontend features multi-vendor support. Next, HLSFactory will apply a design space sampling to reduce the number of designs to be synthesized by HLS and implementation tools (e.g., Vivado), if needed, to shorten the execution time. We will showcase this usage in Section V-C.
Stage ➋ is design synthesis stage, where vendor-specific HLS and implementation tools are invoked to synthesize HLS designs into RTL code and then placed and routed. In this stage, users can submit their HLS designs to be directly synthesized without extrapolating. We will showcase this usage in Section V-D.
Stage ➌ is data aggregation, where statistics and artifacts are collected from the implemented designs and compiled into a tool-agnostic format for use by downstream tasks such as ML training and benchmarking. In this stage, users can submit their synthesized post-implementation results or datasets to be merged. We will showcase this usage in Section V-E.
III-B Stage 1: Design Space Expansion and Sampling
This stage aims at expanding a single HLS design into multiple by enumerating combinations of optimization directives (either inline or in a separate file), such as loop unroll factors, array partitioning schemes, and whether to pipeline loops. Such expansion is critical for ML usage because of two reasons. First, the original HLS designs and benchmarks are far from sufficient for ML training, and obtaining additional HLS designs is challenging. Second, a key application of ML for HLS is to help designers choose the best optimization directives for their HLS designs by predicting post-HLS-synthesis and post-implementation metrics from HLS source code and directives (e.g., [13, 16, 20, 21]). Therefore, an ML-ready HLS dataset must provide wide coverage of how different choices of optimization directives can impact a design. Note that the design space expansion is expected to be across multiple vendors, tools, and devices.
@@ III-B Stage 1: Design Space Expansion and Sampling
III-B1 Vendor-agnostic OptDSL Frontend
@@
Figure 2: A snippet demonstrating the OptDSL syntax.
@@
III-B2 Vendor-specific Concrete Design Generation
@@
The OptDSL file is provided within the abstract design as a file named opt_template.tcl. To lower the abstract design for AMD/Xilinx tools, we generate opt.tcl, a version of opt_template.tcl with bracketed parameters replaced with different concrete values for each design point.
Once these bracketed parameters are substituted, the OptDSL script becomes a valid Tcl script that can be used directly with Vitis HLS.
To support other vendors, the frontend can parse the OptDSL file and identify the specific optimization directives used within it together with their parametrizations. If the provided OptDSL is not sufficient to describe a desired DSE, HLSFactory provides the necessary infrastructure to allow users to specify their own entirely custom frontend as Python code, as long as it conforms to the specified API interface (to be discussed in Sec. IV-A). For instance, a new frontend pass can easily be introduced to parameterize constants in the HLS source code itself: simply copy the existing OptDSL frontend and modify the templating logic and syntax to work with files other than opt_template.tcl.
III-B3 Design Space Sampling
@@
III-C Stage 2: Design Synthesis
@@
Design synthesis is broken down into two steps: (1) HLSSynth, where an HLS design is synthesized to RTL code, and (2) HLSImpl, where the resulting RTL code is implemented, resulting in a fully placed-and-routed design. For AMD/Xilinx designs, Vitis HLS is used for HLSSynth and Vivado for HLSImpl. However, any vendor tool can easily be integrated into the HLSFactory framework, for example, Yosys [22] or Intel HLS (to be demonstrated in Sec. V-C), by providing Python code for the desired ToolFlow subclasses.
III-D Stage 3: Data Extraction and Aggregation
Once all the frontends and tool flows have been executed on a pool of designs, relevant design data must be extracted and aggregated into structured formats.
HLSFactory provides DataAggregator classes to package HLS synthesis data (estimated latency, resource usage), post-implementation data (timing, resource, and power data), tool execution metadata (version, runtime), and build artifacts (LLVM IR, IP blocks) into shareable datasets.
@@ III-D Stage 3: Data Extraction and Aggregation
IV Implementation and Usage
IV-A Vendor Agnostic User API
@@
Figure 3: Example usage of the HLSFactory framework.
@@
TABLE II: The HLSFactory User API.
@@
An example is shown in Fig. 3. The source HLS designs are located and copied to the desired work directory, and the OptDSLFrontend is invoked to sample 10 random design points from each design. The VitisHLSSynthFlow and VitisHLSImplFlow are then be invoked to synthesize and implement each design point, followed by data aggregation using the VitisHLSImplReportFlow to gather data from each implemented design in a standardized format. A full list of the available APIs is available in Table II.
The API also includes abstract base classes (ABCs) that users can subclass to implement their own frontends and tool flows for HLSFactory, for instance, to support another vendor’s HLS tools. HLSFactory abstracts away the complexities of integrating custom user subclasses into the overall dataset generation process, including the use of fine-grained parallelism (to be discussed in Sec. IV-C).
IV-B Directory Structure
Fig. 4 depicts a simple example of the directory structure accepted as input and produced as output of the HLSFactory workflow. As described throughout Sec. III, we first sample the design space for each source abstract design and then run tool flows and data aggregation on the sampled concrete designs. The figure presents the directory structure for the inputs to this process: an abstract design specified in terms of HLS kernel code, an opt_template.tcl file to be used by the OptDSL frontend (described in Sec. III-B1), and auxiliary scripts for the AMD/Xilinx tool flows.
@@ IV-B Directory Structure
Figure 4: The directory structure that HLSFactory uses. Red are input files; green are the intermediate design points; blue are output files.
IV-C Parallel Build Backend
@@
V Evaluations
We evaluate our work through a series of six case studies which demonstrate HLSFactory’s multifunctionality and ease of use.
V-A Case Study 1: Design Space Coverage
We evaluate how the use of design space expansion in HLSFactory quantitatively and qualitatively improves the overall design space of generated datasets in terms of latency (HLS-reported) and resource usage (post-implementation). In the context of ML, improved design space coverage for these metrics is important for robust model training on downstream tasks, such as ML-based QoR prediction. Thus we perform a case study comparing metrics of the base designs in Polybench, MachSuite, and CHStone ($n=29$) with the designs sampled from them ($n=257$).
Figure 5: Effect of design sampling to cover more design space. Sampled designs cover a wider range of metrics than base designs with no optimizations. Latency is HLS estimated; resources are post-implementation. Note that these are stacked density plots to show the effect of cumulative design sampling.

### v3

@@ Abstract.
In addition to its evident application in ML, we showcase the versatility and multi-functionality of our framework through seven case studies: I) Building an ML model for post-implementation QoR prediction; II) Using design space sampling in stage 1 to expand the design space covered from a small base set of HLS designs; III) Demonstrating the speedup from the fine-grained design parallelism backend; IV) Extending HLSFactory to target Intel’s HLS flow across all stages; V) Adding and running new auxiliary designs using HLSFactory; VI) Integration of previously published HLS data in stage 3; VII) Using HLSFactory to perform HLS tool version regression benchmarking.
@@
journalyear: 2024copyright: rightsretainedconference: 2024 ACM/IEEE International Symposium on Machine Learning for CAD; September 9–11, 2024; Salt Lake City, UT, USAbooktitle: 2024 ACM/IEEE International Symposium on Machine Learning for CAD (MLCAD ’24), September 9–11, 2024, Salt Lake City, UT, USAdoi: 10.1145/3670474.3685961isbn: 979-8-4007-0699-8/24/09
[…]
1. Introduction
@@
Machine learning (ML) techniques have been widely applied to different electronic design automation (EDA) flows including high-level synthesis (HLS)  (Mohammadi Makrani et al., 2019; Dai et al., 2018a; Liu and Schafer, 2016; Haaswijk et al., 2018; Luo et al., 2023; Chhabria et al., 2021; Kim et al., 2018; Lin et al., 2020; Singha et al., 2022; Lin et al., 2022) for quality-of-result (QoR) prediction, optimization, and design space exploration (DSE). A key enabler to the success of such ML techniques is high-quality datasets, most of which are developed for individual studies e.g., (Mohammadi Makrani et al., 2019; Singha et al., 2022; Lin et al., 2020). Some recent works have contributed open-source datasets that can be used by other reseachers, e.g.,  (Gautier et al., 2016; Zhou et al., [n. d.]; Wei et al., [n. d.]; Ferretti et al., [n. d.]; Goswami et al., [n. d.]; Bai et al., 2023).
@@
Despite the great benefits of these datasets, there are still fundamental limitations that hinder their wider adoption for ML applications and FPGA research. First, these datasets are usually small or homogeneous, containing only a subset of previously published HLS benchmarks (pol, [n. d.]; Hara et al., [n. d.]; Reagen et al., [n. d.]; Zhou et al., [n. d.]), and frequently consisting exclusively of designs that work with one HLS tool from a single vendor. For example, Spector (Gautier et al., 2016) contains only 9 Intel HLS designs, HLSDataset (Wei et al., [n. d.]) contains 34 AMD/Xilinx HLS designs, and Rosetta (Zhou et al., [n. d.]) contains 6 AMD/Xilinx HLS designs. Second, because of these separately developed HLS datasets, the designs and intermediate/final tool outputs, which serve as important ML model features, are often reported organized in non-standard ad hoc ways. Some datasets contain only source code (Hara et al., [n. d.]; pol, [n. d.]; Reagen et al., [n. d.]), some datasets contain only resource usage and end-to-end throughput (Zhou et al., [n. d.]) but no clock frequency or power numbers, while some contain only post-implementation results (Bai et al., 2023). HLSyn (Bai et al., 2023) is a dataset for HLS designs targeted towards predicting design quality of FPGAs. It consists of a wider range of programs and compiler directives, enabling performance optimization of designs. However, existing datasets require huge manual effort and deep domain-specific knowledge for ML practitioners if they need a complete, unified, and larger dataset, where they must execute all related HLS tools on their own to re-collect and organize the needed information. Third, it is challenging for external users who want to extend the existing datasets by contributing their own designs, primarily caused by ad-hoc data formats and missing details when building these datasets (e.g., tool version, target FPGA device, clock frequency, implementation flow settings). The fundamental limitation is, however, not the lack of another complete and rich HLS dataset, but rather the lack of a flexible and extensible framework to enable continuous contributions to a standardized and sustainable dataset.
@@
  • ML-ready and multi-purpose Beyond simply being used for ML training, as demonstrated with a post-implementation QoR prediction ML model (case study in §5.1), HLSFactory is useful for any task where a large, diverse set of HLS runs is needed, like HLS tool version regression testing (case study in §5.7).
  • High performance and open-source. HLSFactory maximizes parallelism for fast dataset generation of large numbers of designs. HLSFactory is open-source and available on GitHub, including both the end-to-end framework and a large set of sample designs.
In Sec. 2, we first provide background on the prior works in existing HLS benchmarks and datasets. Sec. 3 introduces our HLSFactory framework detailing the three stages. Sec. 4 dives into the implementation of HLSFactory, including how it is configured and extended, and our fine-grained parallelism technique to speed up dataset generation. We then perform several case studies in Sec. 5 that demonstrate the multi-purpose of the proposed framework.
Table 1. A comparison of HLSFactory with the existing work. : feature supported; : feature unsupported; : feature partially supported.
@@
2. Related Work
HLS community has multiple standard benchmarks for assessing HLS tools including PolyBench (pol, [n. d.]), CHStone (Hara et al., [n. d.]), and MachSuite (Reagen et al., [n. d.]), which in total provide around 67 benchmark designs and are far from sufficient for ML training. Rosetta (Zhou et al., [n. d.]), Dai (Dai et al., 2018a), MLSBench (Goswami et al., [n. d.]), DB4HLS (Ferretti et al., [n. d.]), HLSDataset (Wei et al., [n. d.]), and Spector (Gautier et al., 2016) are all recently proposed HLS datasets, where the former four use AMD/Xilinx tools and the last uses Intel tools. MLSBench provides a sampling from different combinations of directives (pragmas) on top of CHStone and MachSuite. DB4HLS provides exhaustive design exploration on 39 designs from MachSuite with a domain-specific language (DSL) for DSE and parallelized synthesis runs. HLSDataset aims to cover all four commonly used benchmarks (PolyBench, CHStone, MachSuite, Rosetta) with a DSL for specifying the design space to sample from. They also illustrate two ML-based case studies for post-implementation resource and power prediction. HLSyn (Bai et al., 2023) uses control data flow graphs (CDFGs) of compiled HLS kernels for QoR prediction using graph neural network approaches; their designs are sampled from PolyBench and MachSuite. The features of selected prior works and HLSFactory are shown in Table 1.
@@ 2. Related Work
3. HLSFactory Framework
[…]
.98Refer to caption
Figure 1. A complete overview of the HLSFactory framework with three stages and three entry points where users can contribute their own designs.
3.1. HLSFactory Overview
@@
Stage ➊ is design space expansion, aiming at expanding a single HLS design into multiple designs by enumerating different combinations of optimization directives (pragmas), which can significantly increase the number of data points for ML applications. In this stage, users can submit one or more HLS designs with possible design space configurations and HLSFactory will extrapolate and expand the complete design space via a frontend. Note that the presented design space expansion stage is explicitly different than traditional design space exploration (commonly abbreviated as DSE). Design space expansion is not optimization guided (as detailed in §3.2.3), so suboptimal designs are included, broadening design space coverage needed to building robust, accurate ML models. This frontend features multi-vendor support and allows for random sampling of generated designs to reduce the number of designs to be synthesized by HLS and implementation tools (e.g., Vivado), if needed for large design spaces, to shorten the execution time.
We will showcase this usage in Section 5.4.
Stage ➋ is design synthesis stage, where vendor-specific HLS and implementation tools are invoked to synthesize HLS designs into RTL code and then placed and routed. In this stage, users can submit their HLS designs to be directly synthesized without extrapolating. We will showcase this usage in Section 5.5.
Stage ➌ is data aggregation, where statistics and artifacts are collected from the implemented designs and compiled into a tool-agnostic format for use by downstream tasks such as ML training and benchmarking. In this stage, users can submit their synthesized post-implementation results or datasets to be merged. We will showcase this usage in Section 5.6.
3.2. Stage 1: Design Space Expansion and Sampling
This stage aims at expanding a single HLS design into multiple by enumerating combinations of optimization directives (either inline or in a separate file), such as loop unroll factors, array partitioning schemes, and whether to pipeline loops. Such expansion is critical for ML usage because of two reasons. First, the original HLS designs and benchmarks are far from sufficient for ML training, and obtaining additional HLS designs is challenging. Second, a key application of ML for HLS is to help designers choose the best optimization directives for their HLS designs by predicting post-HLS-synthesis and post-implementation metrics from HLS source code and directives (e.g., (Wei et al., [n. d.]; Bai et al., 2023; Wu et al., 2021, 2022)). Therefore, an ML-ready HLS dataset must provide wide coverage of how different choices of optimization directives can impact a design. Note that the design space expansion is expected to be across multiple vendors, tools, and devices.
@@ 3.2. Stage 1: Design Space Expansion and Sampling
3.2.1. Vendor-agnostic OptDSL Frontend
@@
Figure 2. A snippet demonstrating the OptDSL syntax.
@@
3.2.2. Vendor-specific Concrete Design Generation
@@
The OptDSL file is provided within the abstract design as a file named opt_template.tcl. To lower the abstract design for AMD/Xilinx tools, we generate opt.tcl, a version of opt_template.tcl with bracketed parameters replaced with different concrete values for each design point. Once these bracketed parameters are substituted, the OptDSL script becomes a valid Tcl script that can be used directly with Vitis HLS.
To support other vendors, the frontend can parse the OptDSL file and identify the specific optimization directives used within it together with their parametrizations. If the provided OptDSL is not sufficient to describe a desired DSE, HLSFactory provides the necessary infrastructure to allow users to specify their own entirely custom frontend as Python code, as long as it conforms to the specified API interface (to be discussed in Sec. 4.1). For instance, a new frontend pass can easily be introduced to parameterize constants in the HLS source code itself: simply copy the existing OptDSL frontend and modify the templating logic and syntax to work with files other than opt_template.tcl.
3.2.3. Design Space Sampling
@@
Unlike design space exploration, HLSFactory’s enumeration and random sampling approaches are not guided or optimization-driven. The focus is solely on collecting a wide range of designs, including suboptimal designs, which are important for building ML datasets and training ML models that can interpolate to as many unseen designs during evaluation and deployment, not just optimal designs.
@@
3.3. Stage 2: Design Synthesis
@@
Design synthesis is broken down into two steps: (1) HLSSynth, where an HLS design is synthesized to RTL code, and (2) HLSImpl, where the resulting RTL code is implemented, resulting in a fully placed-and-routed design. For AMD/Xilinx designs, Vitis HLS is used for HLSSynth and Vivado for HLSImpl. However, any vendor tool can easily be integrated into the HLSFactory framework, for example, Yosys (Wolf and Glaser, 2013) or Intel HLS (to be demonstrated in Sec. 5.4), by providing Python code for the desired ToolFlow subclasses.
3.4. Stage 3: Data Extraction and Aggregation
Once all the frontends and tool flows have been executed on a pool of designs, relevant design data must be extracted and aggregated into structured formats. HLSFactory provides DataAggregator classes to package HLS synthesis data (estimated latency, resource usage), post-implementation data (timing, resource, and power data), tool execution metadata (version, runtime), and build artifacts (LLVM IR, IP blocks) into shareable datasets.
@@ 3.4. Stage 3: Data Extraction and Aggregation
4. Implementation and Usage
4.1. Vendor Agnostic User API
@@
Figure 3. Example usage of the HLSFactory framework.
@@
Table 2. The HLSFactory User API.
@@
An example is shown in Fig. 3. The source HLS designs are located and copied to the desired work directory, and the OptDSLFrontend is invoked to sample 10 random design points from each design. The VitisHLSSynthFlow and VitisHLSImplFlow are then be invoked to synthesize and implement each design point, followed by data aggregation using the VitisHLSImplReportFlow to gather data from each implemented design in a standardized format. A full list of the available APIs is available in Table 2.
The API also includes abstract base classes (ABCs) that users can subclass to implement their own frontends and tool flows for HLSFactory, for instance, to support another vendor’s HLS tools. HLSFactory abstracts away the complexities of integrating custom user subclasses into the overall dataset generation process, including the use of fine-grained parallelism (to be discussed in Sec. 4.3).
4.2. Directory Structure
Fig. 4 depicts a simple example of the directory structure accepted as input and produced as output of the HLSFactory workflow. As described throughout Sec. 3, we first sample the design space for each source abstract design and then run tool flows and data aggregation on the sampled concrete designs. The figure presents the directory structure for the inputs to this process: an abstract design specified in terms of HLS kernel code, an opt_template.tcl file to be used by the OptDSL frontend (described in Sec. 3.2.1), and auxiliary scripts for the AMD/Xilinx tool flows.
@@ 4.2. Directory Structure
[…]
0.95Refer to caption
Figure 4. The directory structure that HLSFactory uses. Red are input files; green are the intermediate design points; blue are output files.
4.3. Parallel Build Backend
@@
5. Evaluations
We evaluate our work through a series of seven case studies which demonstrate HLSFactory’s multifunctionality and ease of use.
5.1. Case Study 1: ML Prediction of Post-Implementation QoR
Figure 5. True-vs-predicted plots for the HLS-based ML QoR model. Test values are shown for models trained on the complete and partial subset of the training design space. “RAE”: Relative Absolute Error ($|\hat{y}-y|/|y-\bar{y}|$), “R2”: Coefficient of Determination
HLS vendor tools provide resource usage estimates (e.g., #LUTs, #FFs, #DSPs, #BRAMs) and timing information (e.g., II violations, clock speed) for designs based on scheduling and binding results. However, HLS-estimated results often deviate significantly from post-implementation resource usage and may not correlate well with critical timing metrics (e.g., worst negative slack and worst hold slack). Previous works, such as S. Dai et al. (Dai et al., 2018b), address this issue by using ML-based models to predict post-implementation quality-of-results (QoR) metrics based on HLS-reported metrics.
We demonstrate that HLSFactory can replicate the approach used by S. Dai et al. (Dai et al., 2018b) to build ML models for post-implementation QoR predictions targeting Vitis HLS and Vivado. We use HLSFactory built-in Polybench, MachSuite, and CHStone design datasets which provide $n=29$ base designs; using the the OptDSL frontend, design space expansion is performed resulting in $n=257$ final designs. HLSFactory’s APIs are also used run tool synthesis and implementation as well as bundle the HLS post-implementation data into a tabular dataset. A histogram-based gradient boosting regression model is then trained to predict post-implementation reported resources and timing metrics using HLS-reported resources, latency, clock speed, and arithmetic/logic operation counts as model inputs. We train our model on an 80%/20% train-test split, as well as a 25% subset of the training data to demonstrate the utility of design space expansion in improving ML model performance.
Our results, shown in Fig. 5, indicate that the $R^{2}$ value and mean relative error are better for the larger training set achieved through design space expansion. We highlight that generating more data points using HLSFactory’s design space expansion will result in higher prediction accuracy, even when randomly sampling from the entire design space and including suboptimal, i.e. “bad”, designs (in terms of QoR metrics). For most resource prediction targets, our ML model also has a lower relative error than the HLS-reported values, showing improvement over the HLS tool itself. These results highlight the utility of HLSFactory applied to ML for EDA and the importance of design space expansion, even with a smaller sample size, for robust ML dataset construction and model training.
5.2. Case Study 2: Design Space Coverage
We evaluate how the use of design space expansion in HLSFactory quantitatively and qualitatively improves the overall design space of generated datasets in terms of latency (HLS-reported) and resource usage (post-implementation). In the context of ML, improved design space coverage for these metrics is important for robust model training on downstream tasks, such as ML-based QoR prediction. Thus we perform a case study comparing metrics of the base designs in Polybench, MachSuite, and CHStone ($n=29$) with the designs sampled from them ($n=257$); this is the same dataset used in §5.1.